Rank each of the following value in order of smallest to largest:

a)The upper quartile

b)The median

c)The 20th percentile

d)Decile 6

e)The 0.38 fractile

f)The 70th percentile

g)The lower quartile


1
Expert's answer
2021-05-10T18:20:44-0400

Let number of observations = N

Then,

First we calculate all of these and then arrange accordingly

So,

Upper Quartile "Q_3=\dfrac{3}{4}N=0.75N"


Median "Q_2=\dfrac{1}{2}N=0.5N"


20th percentile "P_{20}=\dfrac{20}{100}N=0.2N"


Decile 6, "D_6=\dfrac{6}{10}N=0.6N"


0.38 fractile, "F_{0.38}=0.38N"


70th percentile "P_{70}=\dfrac{70}{100}N=0.70N"


Lower Quartile, "Q_1=\dfrac{1}{4}N=0.25N"


From the above results, now arrange from smallest to largest

(1) "20^{th}" Percentile

(2) Lower Quartile

(3) 0.38 fractile

(4) Median

(5) Decile 6

(6) "70^{th}" Percentile

(7) Upper Quartile